OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x − 3y − z if x = −3, y = 2, and z = −2

OpenStudy (ajprincess):

Just plug in the values. That is 2x-3y-z=2(-3)-3(2)-(-2)

OpenStudy (anonymous):

how do you do that ?

OpenStudy (ajprincess):

2x-3y-z =2(-3)-3(2)-(-2) =-6-6+2 =-12+2 =-10 Is it clear?

OpenStudy (anonymous):

okay i see it! thanks (:

OpenStudy (ajprincess):

yw:)